# Rethinking Approximate Gaussian Inference in Classification

### v3

@@ A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula [owen_table_1980, Eq. 10,010.8].
@@ A.3 Probit Approximation for Gaussian Sigmoid Integral
as $y\to 0$. Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Eq.˜32 we derive the probit approximation [spiegelhalter_sequential_1990, mackay_evidence_1992]
@@ A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [lu_mean-field_2021] is obtained as follows
@@
(1) is the mean field approximation, and (3) uses the probit approximation Eq.˜34. [lu_mean-field_2021] provides two other variants of this approximation with other choices of approximation (2).
@@ Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [daganzo_chapter_1979].
@@ Appendix C Theoretical Analyses
In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses, for instance in the synthetic experiment (Fig.˜1) or in [daunizeau_semi-analytical_2017].
@@ C.1 Informal Analysis of Monte Carlo Approximations
In this section we prove Theorem˜3.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
@@
Let $(X_{n})_{n\geq 1},(Y_{n})_{n\geq 1}$ be sequences of $(0,M)$-valued random variables for some $M>0$. Suppose $\sup_{n}\mathbb{E}[Y_{n}^{-8}]<\infty$ and that $\mathrm{Var}(Y_{n})\to 0$ as $n\to\infty$. Then
$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathrm{Var}(Y_{n})^{1/2})$ (48)
as $n\to\infty$.
@@
where $\eta(X)\in[\mathbb{E}[X],X]\text{ or }[X,\mathbb{E}[X]]$ and $\xi(Y)\in[\mathbb{E}[Y],Y]\text{ or }[Y,\mathbb{E}[Y]]$, using the Lagrange form of the remainder. So by the Cauchy-Schwarz inequality,
@@
$\leq\mathbb{E}\left[\frac{1}{\xi(Y)^{8}}\right]^{1/4}\mathbb{E}[(X-\mathbb{E}[X])^{4}]^{1/4}\mathbb{E}[(Y-\mathbb{E}[Y])^{2}]^{1/2}$
$\quad+\mathbb{E}\left[\frac{\eta(X)^{2}}{\xi(Y)^{6}}\right]^{1/2}\mathbb{E}[(Y-\mathbb{E}[Y])^{4}]^{1/2}.$
@@
$\eta(X_{n})$ $\leq M,$ (51)
$\sup_{n}\mathbb{E}\left[\frac{1}{\xi(Y_{n})^{k}}\right]$ $\leq\sup_{n}\mathbb{E}\left[\frac{1}{\min(Y_{n}^{k},\mathbb{E}[Y_{n}]^{k})}\right]\leq\sup_{n}\mathbb{E}\left[\frac{1}{Y_{n}^{k}}+\frac{1}{\mathbb{E}[Y_{n}]^{k}}\right]<\infty\text{ for }k\in\{6,8\},$
$\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])^{4}]$ $\leq M^{4},$
$\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{4}]$ $\leq M^{2}\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}].$
@@
$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}]^{1/2})=O(\mathrm{Var}(Y_{n})^{1/2})$ (52)
as $n\to\infty$. ∎
@@
where $\bm{\Sigma}$ has $\bm{\sigma}$ as diagonal. We now restate Theorem˜3.1 with an explicit expression for $M(\mathcal{K})$:
@@
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (55)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$.
@@
Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, and the proof strategy fails in that case. $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$. By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$. Thus we can apply Lemma˜C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}$, to get
$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)$ (58)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$. Thus Taylor expanding each term of (1) around $p_{c}$ we get
@@
$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)^{2}$
$=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$, where we used that $p_{c}$, $\tilde{p}_{c}$, and hence $\omega_{c}(\tilde{p}_{c})$ is bounded away from 0 by compactness of $\mathcal{K}$.
@@ Appendix D Moment Matching Beta Distributions
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Eq.˜61, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
@@ D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.˜D.1. This extends the work from [mackay_choice_1998], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
@@
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [ay_information_2017, Section 5.1.4]. Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane ay_information_2017. We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [reynolds_hyperbolic_1993],
@@
For $\mathcal{B}((0,1))$, we use the isometric embedding from le_brigant_fisher-rao_2021:
@@
Finally, we choose our embedding Eq.˜63 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.˜64 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq.˜61 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.

### v4

@@ A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula [41, Eq. 10,010.8].
@@ A.3 Probit Approximation for Gaussian Sigmoid Integral
as $y\to 0$. Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Eq.˜32 we derive the probit approximation [47, 32]
@@ A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [31] is obtained as follows
@@
(1) is the mean field approximation, and (3) uses the probit approximation Eq.˜34. [31] provides two other variants of this approximation with other choices of approximation (2).
@@ Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [9].
@@ Appendix C Theoretical Analyses
In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses, for instance in the synthetic experiment (Fig.˜1) or in [10].
@@ C.1 Informal Analysis of Monte Carlo Approximations
In this section we extend and prove Theorem˜3.1, and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
@@
Let $\mathcal{Y}$ be a family of $(0,M)$-valued random variables such that $\sup_{Y\in\mathcal{Y}}\mathbb{E}[Y^{-k}]<\infty$ for all $k>0$. Then for $X,Y$ $(0,M)$-valued random variables with $Y\in\mathcal{Y}$ and any $\epsilon>0$,
$\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|=O\left(\mathrm{Var}(X)^{\frac{1}{2}-\epsilon}\mathrm{Var}(Y)^{\frac{1}{2}}+\mathrm{Var}(Y)^{1-\epsilon}\right)$ (48)
as $\mathrm{Var}(Y)\to 0$ or $\mathrm{Var}(X),\mathrm{Var}(Y)\to 0$.
@@
where $\eta(X)\in[\mathbb{E}[X],X]\text{ or }[X,\mathbb{E}[X]]$ and $\xi(Y)\in[\mathbb{E}[Y],Y]\text{ or }[Y,\mathbb{E}[Y]]$, using the Lagrange form of the remainder. So by the multivariate Hölder inequality we have for any $0<\epsilon<1/2$,
@@
$\leq\mathbb{E}\left[\frac{1}{\xi(Y)^{2/\epsilon}}\right]^{\epsilon}\mathbb{E}[(X-\mathbb{E}[X])^{\frac{2}{1-2\epsilon}}]^{\frac{1}{2}-\epsilon}\mathbb{E}[(Y-\mathbb{E}[Y])^{2}]^{\frac{1}{2}}$
$\quad+\mathbb{E}\left[\frac{\eta(X)^{1/\epsilon}}{\xi(Y)^{3/\epsilon}}\right]^{\epsilon}\mathbb{E}[(Y-\mathbb{E}[Y])^{\frac{2}{1-\epsilon}}]^{1-\epsilon}.$
@@
$\mathbb{E}[(X-\mathbb{E}[X])^{\frac{2}{1-2\epsilon}}]$ $\leq M^{\frac{4\epsilon}{1-2\epsilon}}\mathbb{E}[(X-\mathbb{E}[X])^{2}]=M^{\frac{4\epsilon}{1-2\epsilon}}\mathrm{Var}(X),$ (51)
$\eta(X)^{1/\epsilon}$ $\leq M^{1/\epsilon},$
$\sup_{Y\in\mathcal{Y}}\mathbb{E}\left[\frac{1}{\xi(Y)^{k}}\right]\leq\sup_{Y\in\mathcal{Y}}\mathbb{E}$ $\left[\frac{1}{\min(Y,\mathbb{E}[Y]^{k})}\right]\leq\sup_{Y\in\mathcal{Y}}\mathbb{E}\left[\frac{1}{Y^{k}}+\frac{1}{\mathbb{E}[Y]^{k}}\right]<\infty\text{ for }k\in\{2/\epsilon,3/\epsilon\}.$
@@
$\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|=O\left(\mathrm{Var}(X)^{\frac{1}{2}-\epsilon}\mathrm{Var}(Y)^{\frac{1}{2}}+\mathrm{Var}(Y)^{1-\epsilon}\right)$ (52)
as $\mathrm{Var}(Y)\to 0$ or $\mathrm{Var}(X),\mathrm{Var}(Y)\to 0$. ∎
@@
where $\bm{\Sigma}$ has $\bm{\sigma}$ as diagonal. We now restate Theorem˜3.1 with an explicit expression for $M(\mathcal{K})$ and an additional decay result for the $O$ term:
@@
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O$ (55)
where
$O=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)\qquad\text{as }\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ (56)
and
$O=O\left(\max_{1\leq c\leq C}\mathrm{Var}(Q_{c})^{2-\epsilon}\right)\;\forall\>\epsilon>0\qquad\text{as }\mathrm{Var}(Q_{c})\to 0\;\forall\>1\leq c\leq C.$ (57)
@@
Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-k}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-k}\right]=\infty$, and the proof strategy fails in that case. $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $k>0$. By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma}^{2})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-k}\right]<\infty$ for any $k>0$. Thus we can apply Lemma˜C.1 with $\mathcal{Y}=\left\{\sum_{c=1}^{C}Q_{c}:(\bm{\mu},\bm{\sigma}^{2})\in\mathcal{K}^{C}\right\}$ to get that for any $\epsilon>0$,
$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{1}{2}}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1-\epsilon}\right)$ (60)
as $\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)\to 0$ or $\mathrm{Var}(Q_{c}),\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)\to 0$. Thus Taylor expanding each term of (1) around $p_{c}$ we get
@@
$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\Bigg(\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{1}{2}}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1-\epsilon}\Bigg)^{2}$
$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\Bigg(\mathrm{Var}(Q_{c})^{1-2\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)$
$\quad+\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{3}{2}-\epsilon}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{2-2\epsilon}\Bigg)$
$=O\Bigg(\sum_{c=1}^{C}\mathrm{Var}(Q_{c})^{1-2\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)$
$\quad+\sum_{c=1}^{C}\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{3}{2}-\epsilon}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{2-2\epsilon}\Bigg)$
where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$, where we used that $p_{c}$, $\tilde{p}_{c}$, and hence $\omega_{c}(\tilde{p}_{c})$ is bounded away from 0 by compactness of $\mathcal{K}$. We now distinguish two cases. If we only assume $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, then we see from Eq.˜61 that
$(1)=O\Bigg(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\Bigg).$ (62)
If we make the stronger assumption $\mathrm{Var}(Q_{c})\to 0$ for all $c$, then using the inequality
$\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)=\sum_{1\leq c_{1},c_{2}\leq C}\operatorname{Cov}(Q_{c_{1}},Q_{c_{2}})\leq\sum_{1\leq c_{1},c_{2}\leq C}\mathrm{Var}(Q_{c_{1}})^{1/2}\mathrm{Var}(Q_{c_{2}})^{1/2}$ (63)
we see that $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)=O\left(\max_{1\leq c\leq C}\mathrm{Var}(Q_{c})\right)$ and hence, from Eq.˜61,
$(1)=O\left(\max_{1\leq c\leq C}\mathrm{Var}(Q_{c})^{2-2\epsilon}\right)$ (64)
and the result follows since $\epsilon$ is arbitrary. ∎
@@
Equation˜63 shows that requiring $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ is weaker than requiring $\mathrm{Var}(Q_{c})\to 0$ for all $c$; but we see in Theorem˜C.2 that assuming the latter gives almost twice as strong a decay rate in $\operatorname{D_{KL}}$. This means that neither Eq.˜56 nor Eq.˜57 can be proven from the other; both results are thus of interest.
@@ Appendix D Moment Matching Beta Distributions
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Eq.˜66, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
@@ D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.˜D.1. This extends the work from [33], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting. Also see the discussion in [8].
@@
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [3, Section 5.1.4]. Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al. [3, Example 3.1]. We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [43],
@@
For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. [26, Proposition 2]:
@@
Finally, we choose our embedding Eq.˜68 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.˜69 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq.˜66 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.